OpenStudy (anonymous):

Express each ratio as a unit rate. 140 miles on 6 gallons.

OpenStudy (cwrw238):

140 / 6

OpenStudy (anonymous):

Convert each rate using dimensional analysis. 30 mi/h= ___ft/min

OpenStudy (anonymous):

I don't understand

OpenStudy (texaschic101):

140/6 = 23.33 per gallon 30 mi/hr = ___ ft/min 1 mile = 5280 ft, so 30 miles = (30 * 5280) = 158400 1 hr = 60 min 158400/60 reduces to 2640 ft/min

OpenStudy (anonymous):

Thank you so much I am truly struggling

OpenStudy (texaschic101):

it can get confusing sometimes. I wish you the best of luck :)

OpenStudy (anonymous):

__lb to 14kg a.28.4 b.30.8 c.26.8 d.20.4